# A preorder on the set of links with applications to symmetric unions

### v2

@@ 5. Small links and finiteness
Let us assume that $L$ is not a small knot. By the Claim 5.3, $E(L)$ contains a properly embedded essential surface $S$ whose boundary is empty or a collection of meridian curves and which does not admit any meridional compressing annulus. Let $F\subset\mathcal{O}(L)$ be the closed embedded $2$-suborbifold obtained by gluing each boundary component of $S$ to the corresponding meridian disk with a singular point of branching index $2$. Suppose, to the contrary, that $F$ is compressible in $\mathcal{O}(L)$, that is, there is a compressing disk $\Delta$ for $F$. It must be a discal $2$-orbifold with a singular point of branching index $2$ since $S=F\cap E(L)$ is essential in $E(L)$. But such a compressing disk $\Delta$ corresponds to a meridional compressing annulus $A=\Delta\cap E(L)$ for $S$, which is impossible by the choice of $S$. Hence, the orbifold $\mathcal{O}(L)$ contains an essential closed orientable $2$-suborbifold and so is not small. ∎
@@
Let $L$ be a small link and $L^{\prime}$ a link such that $L\succeq L^{\prime}$. Suppose, to the contrary, that the link $L^{\prime}$ is not small. By Lemma 5.2, the $3$-orbifold $\mathcal{O}(L^{\prime})$ is not small, that is, it contains an orientable closed incompressible $2$-suborbifold $F$. Therefore, its orbifold fundamental group $G^{\mathrm{orb}}(L^{\prime})$ splits along the orbifold fundamental group $\pi_{1}^{\mathrm{orb}}(F)$ as an amalgamated free product or an HNN extension. In particular, $G^{\mathrm{orb}}(L^{\prime})$ acts non-trivially on the Bass-Serre tree $\mathcal{T}$, without edge inversions, associated to this algebraic decomposition. The epimorphism $\varphi\colon G^{\mathrm{orb}}(L)\twoheadrightarrow G^{\mathrm{orb}}(L^{\prime})$ induces a non-trivial action of the group $G^{\mathrm{orb}}(L)$ on the Bass-Serre tree $\mathcal{T}$ without edge inversions. It follows from [64, Corollary 10.2] that the good orbifold $\mathcal{O}(L)$ contains an orientable incompressible $2$-suborbifold, and thus it is not small. By Lemma 5.2, this contradicts the assumption that $L$ is a small link. ∎
@@
(1) A Montesinos knot with three rational tangles is a small knot by [53, Corollary  4.(a)], hence $K^{\prime}$ must be a small knot by Theorem 1.11(1). Since a small knot is prime, then by Theorem 1.5(3) and Remark 1.6, $K^{\prime}$ is the unknot, a $2$-bridge knot, or a Montesinos knot with $r^{\prime}\leq 3$ rational tangles, and now $r^{\prime}=3$.
@@ 5.2. Small links and determinants
Conversely, suppose $\det L=0$. Then, there is an epimorphism from the homology group $H_{1}(\Sigma_{2}(L);\mathbb{Z})$ onto $\mathbb{Z}$. Let $\tau_{\ast}$ and $\tau_{\sharp}$ denote the automorphisms induced by the covering involution $\tau$ of $\Sigma_{2}(L)$ on $\pi_{1}(\Sigma_{2}(L))$ and $H_{1}(\Sigma_{2}(L);\mathbb{Z})$, respectively. By [25, Theorem 1], we have $\tau_{\sharp}=-\mathrm{id}$. Hence, defining an action of $\mathbb{Z}/2\mathbb{Z}$ on $\mathbb{Z}$ by $\mathrm{id}$, we conclude that the epimorphisms $\pi_{1}(\Sigma_{2}(L))\to H_{1}(\Sigma_{2}(L);\mathbb{Z})\to\mathbb{Z}$ are $\mathbb{Z}/2\mathbb{Z}$-equivariant. Therefore, there is an epimorphism
$G^{\mathrm{orb}}(L)\cong\pi_{1}(\Sigma_{2}(L))\rtimes_{\tau_{\ast}}\mathbb{Z}/2\mathbb{Z}\twoheadrightarrow\mathbb{Z}\rtimes_{-\mathrm{id}}\mathbb{Z}/2\mathbb{Z}\cong\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z}.$
Now the proposition follows from the fact that the infinite dihedral group surjects onto every finite dihedral group. ∎
@@
Let $L\subset S^{3}$ be a small link. By Lemma 5.2, the associated $3$-orbifold $\mathcal{O}(L)$ is small. It follows from $\mathit{PSL}(2,\mathbb{C})$-Culler-Shalen theory in the setting of $3$-orbifolds that the $\mathit{PSL}(2,\mathbb{C})$-character variety $\mathcal{X}(\mathcal{O}(L))$ is 0-dimensional, and thus consists of finitely many points. See [22], [15, Sections  3 and 4], and [64, Corollary 10.2].
Suppose, to the contrary, that $\det L=0$. By the proof of Proposition 5.5, $G^{\mathrm{orb}}(L)$ surjects onto the infinite dihedral group $\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z}$. This induces an injection $\mathcal{X}(\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z})\to\mathcal{X}(\mathcal{O}(L))$ between $\mathit{PSL}(2,\mathbb{C})$-character varieties. Here, $\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z}$ surjects onto all finite dihedral groups $D_{2n}$ and those can be realized as subgroups of $\mathit{PSL}(2,\mathbb{C})$ by [45, C.6 and Theorem C.9 in Chapter V] (see also [55, Lemma 4.7]). For $n\geq 2$, the elliptic elements of order $n$ generating the normal subgroup $\mathbb{Z}/n\mathbb{Z}\subset D_{2n}$ have distinct traces. Therefore, $\mathcal{X}(\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z})$ is infinite, which contradicts the fact that $\mathcal{X}(\mathcal{O}(L))$ contains of only finitely many points. ∎
@@
A split link $L=L_{1}\sqcup L_{2}$ such that $L_{2}$ has $\geq 2$ bridges $\pi_{1}$-dominates infinitely many links with $\geq 3$ bridges. This follows from the fact that
@@
Such a pair $(S^{3},L)$ with $L=L_{1}\sqcup L_{2}$ is a good pair in the sense of Kawauchi. It follows from Kawauchi’s almost identical imitation theory that the pair $(S^{3},L)$ admits almost identical imitations $(S^{3},L^{\ast})$ which are hyperbolic links with the same number of components as $L$ (see [36, Theorem 1.1]). By the construction, the imitation map $q\colon(S^{3},L^{\ast})\to(S^{3},L)$ has the property that the restriction $q|\colon(S^{3},L^{\ast}\setminus{K^{\ast}})\to(S^{3},L\setminus{K})$ is homotopic to a diffeomorphism for any connected components $K^{\ast}$ of $L^{\ast}$ with $q(K^{\ast})=K$. Hence the imitation map induces a proper degree-one map from the exterior $E(L^{\ast})$ onto $E(L)$ which sends the meridians of $L^{\ast}$ to those of $L$, and thus induces an epimorphism $G^{\mathrm{orb}}(L^{\ast})\twoheadrightarrow G^{\mathrm{orb}}(L)$. It follows that $L^{\ast}\succeq L=L_{1}\sqcup L_{2}\succeq L(\frac{p}{q})\sharp L_{2}$ for every $2$-bridge link $L(\frac{p}{q})$.
@@
Let $L\subset S^{3}$ be a small link. By Lemma 5.2, the associated $3$-orbifold $\mathcal{O}(L)$ is small. It follows from $\mathit{PSL}(2,\mathbb{C})$-Culler-Shalen theory [15, Section  4] and [64, Corollary 10.2] that the $\mathit{PSL}(2,\mathbb{C})$-character variety $\mathcal{X}(\mathcal{O}(L))$ is 0-dimensional, and thus consists of finitely many points. By Theorem 1.11(1), any link $L^{\prime}$ such that $L\succeq L^{\prime}$ is small and the orbifold $\mathcal{O}(L^{\prime})$ is small by Lemma 5.2. It follows that the orbifold $\mathcal{O}(L^{\prime})$ is irreducible and atoroidal. Hence, by the geometrization theorem of $3$-orbifolds (see [9, Theorem 2]), $\mathcal{O}(L)$ is hyperbolic or Seifert fibered with a small base which is either hyperbolic, Euclidean, elliptic or bad. We distinguish three mutually exclusive cases according to the geometry of $\mathcal{O}(L^{\prime})$ and prove the finiteness result for each case.
@@
given by the faithful discrete representation $\pi^{\prime}\colon\pi_{1}(\mathcal{B}^{\prime})\to\mathit{PSL}(2,\mathbb{R})$ corresponding to the holonomy of a hyperbolic structure on the $2$-orbifold base $\mathcal{B}^{\prime}$. Like in the hyperbolic case above, each link $L^{\prime}$ $\pi$-dominated by $L$ gives an irreducible character $\varphi^{*}(\chi_{\rho^{\prime}})\in\mathcal{X}^{\mathrm{irr}}(\mathcal{O}(L))$ which determines the irreducible representation $\rho^{\prime}\circ\varphi\colon G^{\mathrm{orb}}(L)\to\mathit{PSL}(2,\mathbb{C})$ up to conjugation in $\mathit{PSL}(2,\mathbb{C})$ and thus its image $\rho^{\prime}(G^{\mathrm{orb}}(L^{\prime}))=\pi^{\prime}(\pi_{1}(\mathcal{B}^{\prime}))$ up to isomorphism. Since the representation $\pi^{\prime}$ is faithful, the irreducible character $\varphi^{*}(\chi_{\rho^{\prime}})\in\mathcal{X}^{\mathrm{irr}}(\mathcal{O}(L))$ determines the orbifold group $\pi_{1}(\mathcal{B}^{\prime})$ up to isomorphism. A hyperbolic closed $2$-dimensional orbifold is determined up to homeomorphism by its orbifold group (see [44]), therefore the irreducible character $\varphi^{*}(\chi_{\rho^{\prime}})\in\mathcal{X}^{\mathrm{irr}}(\mathcal{O}(L))$ determines the base $\mathcal{B}^{\prime}$ of the Seifert fibered orbifold $\mathcal{O}(L^{\prime})$ up to homeomorphism. Since there are at most finitely many irreducible characters in $\mathcal{X}^{\mathrm{irr}}(\mathcal{O}(L))$, there are at most finitely many possible topological types for the orbifold base $\mathcal{B}^{\prime}$ of the Seifert fibered orbifold $\mathcal{O}(L^{\prime})$. There is an epimorphism $\tilde{\varphi}\colon\pi_{1}(\Sigma_{2}(K))\twoheadrightarrow\pi_{1}(\Sigma_{2}(L^{\prime}))$ and thus the order $|H_{1}(\Sigma_{2}(L^{\prime});\mathbb{Z})|$ of the first homology group of $\Sigma_{2}(L^{\prime})$ divides $|H_{1}(\Sigma_{2}(L);\mathbb{Z})|$. Therefore, the value $\det L^{\prime}=|H_{1}(\Sigma_{2}(L^{\prime});\mathbb{Z})|$ is bounded and the finiteness of the link types for $L^{\prime}$ follows from Lemma 5.11. ∎
@@ 6. $\pi$-minimal links
A link that is not prime is either split or a non-trivial connected sum. A split link $L=L_{1}\sqcup L_{2}$ $\pi$-dominates a connected sum $L_{1}\sharp L_{2}$ and thus it is not minimal unless $L_{1}=U$ is the unknot and $L_{2}=L(\frac{p}{q})$ is a $2$-bridge link. Therefore,
$G^{\mathrm{orb}}(L_{1}\sqcup L_{2})=G^{\mathrm{orb}}(U\sqcup L(\tfrac{p}{q}))=\mathbb{Z}/2\mathbb{Z}\ast D_{2p}=(\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z})\ast_{\mathbb{Z}/2\mathbb{Z}}D_{2p}.$
The free product $\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z}$ surjects onto any dihedral group $D_{2q}$, therefore the group $G^{\mathrm{orb}}(L_{1}\sqcup L_{2})$ surjects onto the group $D_{2q}\ast_{\mathbb{Z}/2\mathbb{Z}}D_{2p}$, for any $q\geq 1$. It follows that $L_{1}\sqcup L_{2}$ cannot be $\pi$-minimal unless $L_{2}$ is the unknot and $L_{1}\sqcup L_{2}$ is the $2$-component trivial link.
Let $L=L_{1}\sharp L_{2}$ with $L_{1}$ and $L_{2}$ not being the unknot. If $L$ is $\pi$-minimal, $L_{1}$ and $L_{2}$ must be $2$-bridge links since $L$ $\pi$-dominates $L_{1}$ and $L_{2}$ and $G^{\mathrm{orb}}(L)$ is not isomorphic to $G^{\mathrm{orb}}(L_{1})$ or to $G^{\mathrm{orb}}(L_{1})$. This follows from the fact that $\pi_{1}(\Sigma_{2}(L))=\pi_{1}(\Sigma_{2}(L_{1}))\ast\pi_{1}(\Sigma_{2}(L_{2}))$ is not isomorphic to $\pi_{1}(\Sigma_{2}(L_{1}))$ or $\pi_{1}(\Sigma_{2}(L_{2}))$ when none of these groups is trivial. ∎
@@
Let $L\subset S^{3}$ be a link with at least three bridges and let $\Sigma_{2}(L)$ be its $2$-fold branched covering. If $\pi_{1}(\Sigma_{2}(L))$ does not surject onto the fundamental group of a closed orientable $3$-manifold which is not cyclic, nor isomorphic to $\pi_{1}(\Sigma_{2}(L))$, then $L$ is $\pi$-minimal.
@@
Let $L\subset S^{3}$ be a link with at least three bridges that is not $\pi$-minimal. Then $L\succeq L^{\prime}$ where $L^{\prime}$ has at least three bridges and there is an epimorphism $\varphi\colon G^{\mathrm{orb}}(L)\twoheadrightarrow G^{\mathrm{orb}}(L^{\prime})$ which is not an isomorphism. This epimorphism induces an epimorphism $\tilde{\varphi}\colon\pi_{1}(\Sigma_{2}(L))\twoheadrightarrow\pi_{1}(\Sigma_{2}(L^{\prime}))$. Since $L^{\prime}$ has at least three bridges, $\Sigma_{2}(L^{\prime})$ cannot be cyclic or trivial, otherwise $L^{\prime}$ would be a $2$-bridge link or the unknot by the orbifold theorem [9, Theorem 1]. If $\pi_{1}(\Sigma_{2}(L^{\prime}))$ is isomorphic to $\pi_{1}(\Sigma_{2}(L))$, then $\tilde{\varphi}$ is an isomorphism because $3$-manifold groups are hopfian by [31]. Then the epimorphism $\varphi\colon G^{\mathrm{orb}}(L)\twoheadrightarrow G^{\mathrm{orb}}(L^{\prime})$ induces an epimorphism $\bar{\varphi}\colon G^{\mathrm{orb}}(L)/\pi_{1}(\Sigma_{2}(L))=\mathbb{Z}/2Z\twoheadrightarrow G^{\mathrm{orb}}(L^{\prime})/\pi_{1}(\Sigma_{2}(L^{\prime}))=\mathbb{Z}/2Z$, since $\varphi(G^{\mathrm{orb}}(L))\not\subset\pi_{1}(\Sigma_{2}(L^{\prime}))$. This epimorphism $\bar{\varphi}$ must be an isomorphism, and hence $\varphi$ itself is an isomorphism which is not possible. Therefore, if $L$ is not $\pi$-minimal and has at least three bridges, then the group $\pi_{1}(\Sigma_{2}(L))$ surjects onto the fundamental group of a closed $3$-manifold which is neither trivial, cyclic, nor isomorphic to $\pi_{1}(\Sigma_{2}(L))$. ∎
@@
Let $K$ be a $(p,q)$-torus knot with $2\leq p<q$ and $p$, $q$ are coprime. Let us assume that $p$ and $q$ are prime. If $p$ or $q$ is even, then $p=2$ and the torus knot $T(2,q)$ is $\pi$-minimal. So, in case (1), we can assume that $p$ and $q$ are odd primes. Then it follows that in both cases (1) and (2) the $2$-fold branched cover $\Sigma_{2}(K)$ is Seifert fibered with orbifold base $S^{2}(\alpha_{1},\alpha_{2},\alpha_{3})$, where $2\leq\alpha_{1}<\alpha_{2}<\alpha_{3}$ are prime numbers. One has $(\alpha_{1},\alpha_{2},\alpha_{3})=(2,p,q)$ in case (1). In case (2), since $K$ is assumed not to be an elliptic Montesinos knot, the hypotheses imply that $\frac{1}{\alpha_{1}}+\frac{1}{\alpha_{2}}+\frac{1}{\alpha_{3}}<1$ unless $(\alpha_{1},\alpha_{2},\alpha_{3})=(2,3,5)$ and $K$ is the $(3,5)$-torus knot. Then the proof of Proposition 6.6 follows from Lemma 6.5 and the next claim. ∎

### v3

@@ 5. Small links and finiteness
Let us assume that $L$ is not a small link. By the Claim 5.3, $E(L)$ contains a properly embedded essential surface $S$ whose boundary is empty or a collection of meridian curves and which does not admit any meridional compressing annulus. Let $F\subset\mathcal{O}(L)$ be the closed embedded $2$-suborbifold obtained by gluing each boundary component of $S$ to the corresponding meridian disk with a singular point of branching index $2$. Suppose, to the contrary, that $F$ is compressible in $\mathcal{O}(L)$, that is, there is a compressing disk $\Delta$ for $F$. It must be a discal $2$-orbifold with a singular point of branching index $2$ since $S=F\cap E(L)$ is essential in $E(L)$. But such a compressing disk $\Delta$ corresponds to a meridional compressing annulus $A=\Delta\cap E(L)$ for $S$, which is impossible by the choice of $S$. Hence, the orbifold $\mathcal{O}(L)$ contains an essential closed orientable $2$-suborbifold and so is not small. ∎
@@
Let $L$ be a small link and $L^{\prime}$ a link such that $L\succeq L^{\prime}$. Suppose, to the contrary, that the link $L^{\prime}$ is not small. By Lemma 5.2, the $3$-orbifold $\mathcal{O}(L^{\prime})$ is not small, that is, it contains an orientable closed incompressible $2$-suborbifold $F$. Therefore, its orbifold fundamental group $G^{\mathrm{orb}}(L^{\prime})$ splits along the orbifold fundamental group $\pi_{1}^{\mathrm{orb}}(F)$ as an amalgamated free product or an HNN extension. In particular, $G^{\mathrm{orb}}(L^{\prime})$ acts non-trivially on the Bass-Serre tree $\mathcal{T}$, without edge inversions, associated to this algebraic decomposition. The epimorphism $\varphi\colon G^{\mathrm{orb}}(L)\twoheadrightarrow G^{\mathrm{orb}}(L^{\prime})$ induces a non-trivial action of the group $G^{\mathrm{orb}}(L)$ on the Bass-Serre tree $\mathcal{T}$ without edge inversions. It follows from [65, Corollary 10.2] that the good orbifold $\mathcal{O}(L)$ contains an orientable incompressible $2$-suborbifold, and thus it is not small. By Lemma 5.2, this contradicts the assumption that $L$ is a small link. ∎
@@
(1) A Montesinos knot with three rational tangles is a small knot by [53, Corollary 4.(a)], hence $K^{\prime}$ must be a small knot by Theorem 1.10(1). Since a small knot is prime, then by Theorem 1.4(3) and Remark 1.5, $K^{\prime}$ is the unknot, a $2$-bridge knot, or a Montesinos knot with $r^{\prime}\leq 3$ rational tangles, and now $r^{\prime}=3$.
@@ 5.2. Small links and determinants
Conversely, suppose $\det L=0$. Then, there is an epimorphism from the homology group $H_{1}(\Sigma_{2}(L);\mathbb{Z})$ onto $\mathbb{Z}$. Let $\tau_{\ast}$ and $\tau_{\sharp}$ denote the automorphisms induced by the covering involution $\tau$ of $\Sigma_{2}(L)$ on $\pi_{1}(\Sigma_{2}(L))$ and $H_{1}(\Sigma_{2}(L);\mathbb{Z})$, respectively. By [25, Theorem 1], we have $\tau_{\sharp}=-\mathrm{id}$. Hence, defining an action of $\mathbb{Z}/2\mathbb{Z}$ on $\mathbb{Z}$ by $\mathrm{id}$, we conclude that the epimorphisms $\pi_{1}(\Sigma_{2}(L))\to H_{1}(\Sigma_{2}(L);\mathbb{Z})\to\mathbb{Z}$ are $\mathbb{Z}/2\mathbb{Z}$-equivariant. Therefore, one gets a commutative diagram
${G^{\mathrm{orb}}(L)\cong\pi_{1}(\Sigma_{2}(L))\rtimes_{\tau_{\ast}}\mathbb{Z}/2\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi_{1}}$${\mathbb{Z}/2\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{id}}$${\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z}\cong\mathbb{Z}\rtimes_{-\mathrm{id}}\mathbb{Z}/2\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{Z}/2\mathbb{Z}.}$
So $L$ $\pi$-dominates the split link $U\sqcup U$, which in turn $\pi$-dominates every $2$-bridge link because of the commutative diagram
${\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z}\cong\mathbb{Z}\rtimes_{-\mathrm{id}}\mathbb{Z}/2\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi_{2}}$${\mathbb{Z}/2\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{id}}$${D_{2n}\cong\mathbb{Z}/n\mathbb{Z}\rtimes_{-\mathrm{id}}\mathbb{Z}/2\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{Z}/2\mathbb{Z}.}$
This completes the proof. ∎
@@
Let $L\subset S^{3}$ be a small link. By Lemma 5.2, the associated $3$-orbifold $\mathcal{O}(L)$ is small. It follows from $\mathit{PSL}(2,\mathbb{C})$-Culler-Shalen theory in the setting of $3$-orbifolds that the $\mathit{PSL}(2,\mathbb{C})$-character variety $\mathcal{X}(\mathcal{O}(L))$ is 0-dimensional, and thus consists of finitely many points. See [22], [15, Sections  3 and 4], and [65, Corollary 10.2].
Suppose, to the contrary, that $\det L=0$. By the proof of Proposition 5.5, $G^{\mathrm{orb}}(L)$ surjects onto the infinite dihedral group $\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z}$. This induces an injection $\mathcal{X}(\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z})\to\mathcal{X}(\mathcal{O}(L))$ between $\mathit{PSL}(2,\mathbb{C})$-character varieties. Here, $\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z}$ surjects onto all finite dihedral groups $D_{2n}$ and those can be realized as subgroups of $\mathit{PSL}(2,\mathbb{C})$ by [45, C.6 and Theorem C.9 in Chapter V] (see also [55, Lemma 4.7]). For $n\geq 2$, the elliptic elements of order $n$ generating the normal subgroup $\mathbb{Z}/n\mathbb{Z}\subset D_{2n}$ have distinct traces. Therefore, $\mathcal{X}(\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z})$ is infinite, which contradicts the fact that $\mathcal{X}(\mathcal{O}(L))$ contains only finitely many points. ∎
@@
A split link $L=L_{1}\sqcup L_{2}$ such that $L_{2}$ has $\geq 2$ bridges $\pi$-dominates infinitely many links with $\geq 3$ bridges. This follows from the fact that
@@
Such a pair $(S^{3},L)$ with $L=L_{1}\sqcup L_{2}$ is a good pair in the sense of Kawauchi. It follows from Kawauchi’s almost identical imitation theory that the pair $(S^{3},L)$ admits almost identical imitations $(S^{3},L^{\ast})$ which are hyperbolic links with the same number of components as $L$ (see [36, Theorem 1.1]). By the construction, the imitation map $q\colon(S^{3},L^{\ast})\to(S^{3},L)$ has the property that the restriction $q|\colon(S^{3},L^{\ast}\setminus{K^{\ast}})\to(S^{3},L\setminus{K})$ is homotopic to a diffeomorphism for any connected components $K^{\ast}$ of $L^{\ast}$ with $q(K^{\ast})=K$. Hence, the imitation map induces a proper degree-one map from the exterior $E(L^{\ast})$ onto $E(L)$ which sends the meridians of $L^{\ast}$ to those of $L$. This degree-one map is $\pi_{1}$-surjective, and thus induces the commutative diagram
${G^{\mathrm{orb}}(L^{\ast})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{Z}/2\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{id}}$${G^{\mathrm{orb}}(L)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{Z}/2\mathbb{Z}.}$
It follows that $L^{\ast}\succeq L=L_{1}\sqcup L_{2}\succeq L(\frac{p}{q})\sharp L_{2}$ for every $2$-bridge link $L(\frac{p}{q})$.
@@
Let $L\subset S^{3}$ be a small link. By Lemma 5.2, the associated $3$-orbifold $\mathcal{O}(L)$ is small. It follows from $\mathit{PSL}(2,\mathbb{C})$-Culler-Shalen theory [15, Section  4] and [65, Corollary 10.2] that the $\mathit{PSL}(2,\mathbb{C})$-character variety $\mathcal{X}(\mathcal{O}(L))$ is 0-dimensional, and thus consists of finitely many points. By Theorem 1.10(1), any link $L^{\prime}$ such that $L\succeq L^{\prime}$ is small and the orbifold $\mathcal{O}(L^{\prime})$ is small by Lemma 5.2. It follows that the orbifold $\mathcal{O}(L^{\prime})$ is irreducible and atoroidal. Hence, by the geometrization theorem of $3$-orbifolds (see [9, Theorem 2]), $\mathcal{O}(L)$ is hyperbolic or Seifert fibered with a small base which is either hyperbolic, Euclidean, elliptic or bad. We distinguish three mutually exclusive cases according to the geometry of $\mathcal{O}(L^{\prime})$ and prove the finiteness result for each case.
@@
given by the faithful discrete representation $\pi^{\prime}\colon\pi_{1}(\mathcal{B}^{\prime})\to\mathit{PSL}(2,\mathbb{R})$ corresponding to the holonomy of a hyperbolic structure on the $2$-orbifold base $\mathcal{B}^{\prime}$. Like in the hyperbolic case above, each link $L^{\prime}$ $\pi$-dominated by $L$ gives an irreducible character $\varphi^{*}(\chi_{\rho^{\prime}})\in\mathcal{X}^{\mathrm{irr}}(\mathcal{O}(L))$ which determines the irreducible representation $\rho^{\prime}\circ\varphi\colon G^{\mathrm{orb}}(L)\to\mathit{PSL}(2,\mathbb{C})$ up to conjugation in $\mathit{PSL}(2,\mathbb{C})$ and thus its image $\rho^{\prime}(G^{\mathrm{orb}}(L^{\prime}))=\pi^{\prime}(\pi_{1}(\mathcal{B}^{\prime}))$ up to isomorphism. Since the representation $\pi^{\prime}$ is faithful, the irreducible character $\varphi^{*}(\chi_{\rho^{\prime}})\in\mathcal{X}^{\mathrm{irr}}(\mathcal{O}(L))$ determines the orbifold group $\pi_{1}(\mathcal{B}^{\prime})$ up to isomorphism. A hyperbolic closed $2$-dimensional orbifold is determined up to homeomorphism by its orbifold fundamental group (see [44]), therefore the irreducible character $\varphi^{*}(\chi_{\rho^{\prime}})\in\mathcal{X}^{\mathrm{irr}}(\mathcal{O}(L))$ determines the base $\mathcal{B}^{\prime}$ of the Seifert fibered orbifold $\mathcal{O}(L^{\prime})$ up to homeomorphism. Since there are at most finitely many irreducible characters in $\mathcal{X}^{\mathrm{irr}}(\mathcal{O}(L))$, there are at most finitely many possible topological types for the orbifold base $\mathcal{B}^{\prime}$ of the Seifert fibered orbifold $\mathcal{O}(L^{\prime})$. There is an epimorphism $\tilde{\varphi}\colon\pi_{1}(\Sigma_{2}(K))\twoheadrightarrow\pi_{1}(\Sigma_{2}(L^{\prime}))$ and thus the order $|H_{1}(\Sigma_{2}(L^{\prime});\mathbb{Z})|$ of the first homology group of $\Sigma_{2}(L^{\prime})$ divides $|H_{1}(\Sigma_{2}(L);\mathbb{Z})|$. Therefore, the value $\det L^{\prime}=|H_{1}(\Sigma_{2}(L^{\prime});\mathbb{Z})|$ is bounded and the finiteness of the link types for $L^{\prime}$ follows from Lemma 5.11. ∎
@@ 6. $\pi$-minimal links
A link that is not prime is either split or a non-trivial connected sum. By Proposition 2.9(4), a split link $L=L_{1}\sqcup L_{2}$ $\pi$-dominates a connected sum $L_{1}\sharp L_{2}$, and thus it is not minimal unless $L_{1}=U$ is the unknot and $L_{2}$ is a $2$-bridge link. Therefore, $L=U\sqcup L_{2}=(U\sqcup U)\sharp L_{2}$. By the proof of Proposition 5.5, the split link $U\sqcup U$ $\pi$-dominates every $2$-bridge link, so $L=U\sqcup L_{2}$ $\pi$-dominates the connected sum of $L_{2}$ with every $2$-bridge link. It follows that $L=U\sqcup L_{2}$ cannot be $\pi$-minimal unless $L_{2}$ is the unknot and $L=U\sqcup U$ is the $2$-component trivial link.
Let $L=L_{1}\sharp L_{2}$ with $L_{1}$ and $L_{2}$ not being the unknot. If $L$ is $\pi$-minimal, $L_{1}$ and $L_{2}$ must be $2$-bridge links since $L$ $\pi$-dominates $L_{1}$ and $L_{2}$ by Proposition 2.9(3), and $G^{\mathrm{orb}}(L)$ is not isomorphic to $G^{\mathrm{orb}}(L_{1})$ or to $G^{\mathrm{orb}}(L_{1})$. This follows from the fact that $\pi_{1}(\Sigma_{2}(L))=\pi_{1}(\Sigma_{2}(L_{1}))\ast\pi_{1}(\Sigma_{2}(L_{2}))$ is not isomorphic to $\pi_{1}(\Sigma_{2}(L_{1}))$ or $\pi_{1}(\Sigma_{2}(L_{2}))$ when none of these groups is trivial. ∎
@@
Let $L\subset S^{3}$ be a link with at least three bridges. If $\pi_{1}(\Sigma_{2}(L))$ does not surject onto the fundamental group of a closed orientable $3$-manifold which is neither trivial, cyclic, nor isomorphic to $\pi_{1}(\Sigma_{2}(L))$, then $L$ is $\pi$-minimal.
@@
Let $L\subset S^{3}$ be a link with at least three bridges that is not $\pi$-minimal. Then $L\succeq L^{\prime}$, where $L^{\prime}$ has at least three bridges and there is an epimorphism $\varphi\colon G^{\mathrm{orb}}(L)\twoheadrightarrow G^{\mathrm{orb}}(L^{\prime})$ which is not an isomorphism. By Lemma 2.10, this epimorphism induces an epimorphism $\tilde{\varphi}\colon\pi_{1}(\Sigma_{2}(L))\twoheadrightarrow\pi_{1}(\Sigma_{2}(L^{\prime}))$. Since $L^{\prime}$ has at least three bridges, $\Sigma_{2}(L^{\prime})$ cannot be cyclic or trivial, otherwise $L^{\prime}$ would be a $2$-bridge link or the unknot by the orbifold theorem [9, Theorem 1]. Suppose, to the contrary, that $\pi_{1}(\Sigma_{2}(L^{\prime}))$ is isomorphic to $\pi_{1}(\Sigma_{2}(L))$. Then, $\tilde{\varphi}$ is an isomorphism because $3$-manifold groups are hopfian by [31]. Hence, the commutative diagram below (see also the proof of Lemma 2.10) shows that $\varphi$ itself is an isomorphism, which is not possible:
[…]
Therefore, the group $\pi_{1}(\Sigma_{2}(L))$ surjects onto the fundamental group of a closed $3$-manifold which is neither trivial, cyclic, nor isomorphic to $\pi_{1}(\Sigma_{2}(L))$. ∎
@@
Let $K$ be a $(p,q)$-torus knot with $2\leq p<q$ and $p$, $q$ coprime. Let us assume that $p$ and $q$ are prime. If $p$ or $q$ is even, then $p=2$ and the torus knot $T(2,q)$ is $\pi$-minimal. So, in case (1), we can assume that $p$ and $q$ are odd primes. Then it follows that in both cases (1) and (2) the $2$-fold branched cover $\Sigma_{2}(K)$ is Seifert fibered with orbifold base $S^{2}(\alpha_{1},\alpha_{2},\alpha_{3})$, where $2\leq\alpha_{1}<\alpha_{2}<\alpha_{3}$ are prime numbers. One has $(\alpha_{1},\alpha_{2},\alpha_{3})=(2,p,q)$ in case (1). In case (2), since $K$ is assumed not to be an elliptic Montesinos knot, the hypotheses imply that $\frac{1}{\alpha_{1}}+\frac{1}{\alpha_{2}}+\frac{1}{\alpha_{3}}<1$ unless $(\alpha_{1},\alpha_{2},\alpha_{3})=(2,3,5)$ and $K$ is the $(3,5)$-torus knot. Then the proof of Proposition 6.6 follows from Lemma 6.5 and the next claim. ∎